OpenStudy (anonymous):

Mr. Smith had $2000 to invest. He invested part of it at 5 percent per year and the remainder at 4 percent per year. After one year, his investment grew to $2095. How much of the original investment was at the 5 percent rate?

OpenStudy (anonymous):

viktor!!!!

OpenStudy (anonymous):

Let A=invested at 5% and B=invested at 4% so A+B=2000---- (i) also total returns=$95= 5% of A + 4% of B or 95=0.05A+0.04B----(ii) multiplying (i) by 0.04 we have 0.04A +0.04B = 80 subtracting this from (ii), 15=0.01A or A=1500-->B=500